OpenStudy (anonymous):

3. A ball is thrown horizontally from the roof of a building 55 m tall with a speed of 3.8 m/s. a. How much later does it hit the ground? b. How far from the building will it land? c. What is the velocity of the ball just before it hits the ground?

OpenStudy (anonymous):

Can anyone help! I am really stuck.

OpenStudy (anonymous):

So we can use what we know to solve for the time it takes to hit the ground. Change in height is -55 meters. And initial velocity in y- direction is 0. And acceleration in y-direction is of course -9.8 m/s^2. Does that give you a good head start?

OpenStudy (anonymous):

Use the kinematic equations

OpenStudy (shamim):

a) u hv to use the frmula h=ut+(1/2)*g*t^2 Given u=0 h=55m g=9.8m/s^2 t=?

OpenStudy (shamim):

I wanna help u b, c when u will b online